In a study of speed dating, male subjects were asked to rate the attractiveness of their female dates, and a sample of the results is listed below (1 = not attractive; 10 = extremely attractive). Construct a confidence interval using a 95% confidence level. What do the results tell about the mean attractiveness ratings of the population of all adult females?

8 , 8 , 1 , 8 , 6 , 5 , 6 , 8 , 9 , 9 , 4 , 9

What is the confidence interval for the population mean u?

____ < u > ____

Thanks in advance!
Answer #1

Sample size = n = 12

Sample mean = \bar{x} = 6.75

Standard deviation = s = 2.4541

We have to construct 95% confidence interval.

Formula is

(\bar{x}-E , \bar{x}+E)

Here E is a margin of error.

E=

Degrees of freedom = n - 1 = 12 - 1 = 11

Level of significance = 0.05

tc = 2.201 ( Using t table)

2.20i * 4541 E, = = 1.5593 V12 2

So confidence interval is ( 6.75 - 1.5593 , 6.75 + 1.5593) = > ( 5.1907 , 8.3093)